how do you solve 3+2/5b=11-2/5b?
Answer by stanbon(75887)

solve 3 + (2/5)b = 11 - (2/5)b
---
(4/5)b = 8
---
b = (5/4)8
----
b = 10
=========
Cheers,
Stan H.
===========